OpenStudy (anonymous):

Decide which of the following are solutions to this system of equations: y = x - 5 and y = 3x^2 - 5 (a) (2,7) (b) (0,-5) (c) (3,-2) (d) (2,-3) Both a and c. Both b and d. Only b. Only d.

OpenStudy (solomonzelman):

I bet you don't know what your second equation is? IS it a circle, ellipse hyperbola parabola or a line.... I always see you ask identical questions and apparently without moving on... I won't help you, sorry -:(

OpenStudy (solomonzelman):

If you show me the center of the second equation and tell me what it is, I'll help you.

OpenStudy (anonymous):

It's a line right? and they aren't identical. I got C for my answer

OpenStudy (solomonzelman):

y=x-5 is a line, how about ` y=3x²-5 ` ?

OpenStudy (solomonzelman):

do not look at the line, because all of points in the answer choices are on the line.

OpenStudy (anonymous):

Proabola

OpenStudy (solomonzelman):

Yes

OpenStudy (solomonzelman):

do not look at the line, because all of points in the answer choices are on the line.

OpenStudy (camerondoherty):

*Parabola

OpenStudy (anonymous):

Yes sorry. hah

OpenStudy (solomonzelman):

y=3x²-5 y=3(x-0)²-5 y=3(x-0)²+(-5) → y=a(x-h)²+k where (h,k) is the center.

OpenStudy (solomonzelman):

They gave you (luckily) the point on the line intersecting with the center of the parabola.

OpenStudy (anonymous):

Yes but is it C?